Juanita has 22000 to invest in two accounts that pay simple interest on an annual basis. one account pays 3% in simple interest, and other pays 6% simple interest how much would she have to invest in each account to earn a total interest of $810 after one year?
Answer by mananth(16946)

Part I 3.00% per annum ------------- Amount invested =x
Part II 6.00% per annum ------------ Amount invested = y
22000
Interest----- 810.00

Part I 3.00% per annum ---x
Part II 6.00% per annum ---y
Total investment
x + 1 y= 22000 -------------1
Interest on both investments
3.00% x + 6.00% y= 810
Multiply by 100
3 x + 6 y= 81000.00 --------2
Multiply (1) by -3
we get
-3 x -3 y= -66000.00
Add this to (2)
0 x 3 y= 15000
divide by 3
y = 5000
Part I 3.00% $ 17000
Part II 6.00% $ 5000

CHECK
17000 --------- 3.00% ------- 510.00
5000 ------------- 6.00% ------- 300.00
Total -------------------- 810.00

m.ananth@hotmail.ca